OpenStudy (anonymous):

A rectangle has a perimeter of 26 feet. Twice its length (l) is one foot more than three times its width (w). What is the length of the rectangle, in feet? A. 5 B. 13 C. 8 D. 16

OpenStudy (anonymous):

it is c 2l-1=3w

OpenStudy (anonymous):

i had to do some guess and check

OpenStudy (whpalmer4):

shouldn't be necessary to guess. \[ 2l + 2w = 26\](perimeter is 26 feet)\[ 2l = 3w+1 \](twice its length is 1 foot more than thrice its width) Substitute the second equation into the first:\[3w+1+2w=26\]Collect like terms:\[5w=25\]\[w=5\]\[2l=3w+1\]\[2l=3(5)+1=16\]\[l=16/2=8\]